Find the probabilities of the following.

1. 𝑃(𝑧 > 1.36)

2. 𝑃(𝑧 < 2.45)

3. 𝑃(1.2 < 𝑧 < 1.4)

4. 𝑃(βˆ’2.75 < 𝑧 < βˆ’0.56)

5. 𝑃(𝑧 > βˆ’1.05)


1
Expert's answer
2021-03-23T14:45:13-0400

1. "P(z>1.36)=1-P(z\leq 1.36)"



"\approx1-0.913085\approx0.086915"

2. "P(z<2.45)\approx0.992857"


3. "P(1.2<z<1.4)=P(z<1.4)-P(z\leq 1.2)"



"\approx0.9192433-0.8849303\approx0.034313"

4. "P(-2.75<z<-0.56)=P(z<-0.56)-P(z\leq -2.75)"



"\approx0.2877397-0.0029798\approx0.284760"

5. "P(z>-1.05)=1-P(z\leq -1.05)"



"\approx1-0.146859\approx0.853141"